Sharp Discounts Wholesale Club has two service desks, one at each entrance of the store. Customers arrive according to a Poisson distribution at each entrance at an average of one every 7.5 minutes. Without exception, when customers finish shopping, they checkout at the service desk near the entrance where they entered the store. The negative exponentially distributed service time is 6 minutes per customer.

How many customers, on average, are in each service desk system (waiting and being served)?
Answer #1

IN this case, there are two entry - service systems, where a person entering from one location would get billed near the desk at the same entrance.

Arrival rate at each desk lambda = 60 / 7.5 =8 per hour

service rate mu= 60/6 =10 per hour

Number of customers in each service - desk system = lambda / mu-lambda = 8/10-8 = 4